OpenStudy (anonymous):

An airplane is flying at an altitude of 7 mi and passes directly over a radar antenna. When the plane is 10 mi from the antenna (s=10), the radar detects that the distance s is changing at the rate 300 mph. What is the speed of the airplane at that moment?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

Is this a Doppler effect problem?

OpenStudy (anonymous):

no, it's a related rates problem

OpenStudy (anonymous):



OpenStudy (anonymous):

this look right so far?

OpenStudy (anonymous):

lol yea, mine is close 

OpenStudy (anonymous):

Okay, so I'm going to assume that the radar is actually measuring \(s'\) and what we really want is \(x'\).

OpenStudy (anonymous):

well... i know that I want to find dx/dt right?

OpenStudy (anonymous):

I'm also going to assume it's not changing altitude \((y'=0)\) and that \(x'\) is constant.

OpenStudy (anonymous):

so is 7 constant right?

OpenStudy (anonymous):

and I would do the derivative of Pythagorean theorem equation?

OpenStudy (anonymous):

We have \[ \begin{array}{rcl} s^2 &=& y^2 + x^2 \\ 2ss' &=& 2yy' + 2xx' \end{array}\]

OpenStudy (anonymous):

This is just the chain rule.

OpenStudy (anonymous):

but what you did looks similar to Pythagorean lol

OpenStudy (anonymous):

Yeah there is Pythagorean in there.

OpenStudy (anonymous):

yea i got the same thing... and then i stopped lol

OpenStudy (anonymous):

Well you know \(s=10, s'=300, y=7,y'=0,x=\sqrt{10^2-7^2} \)

OpenStudy (anonymous):

What do you get when you plug that in and solve for \(x'\)? Is it correct?

OpenStudy (anonymous):

Wait, hold on, I think I might have missed something.

OpenStudy (anonymous):

Are you getting the correct answer?

OpenStudy (anonymous):

well if y is constant, wouldn't there be no dy/dx * y

OpenStudy (anonymous):

i got square root of 51

OpenStudy (anonymous):

Yeah, it would essentially be 0.

OpenStudy (anonymous):

for x

OpenStudy (anonymous):

yeah, but for \(x'\)?

OpenStudy (anonymous):

isn't that what we're trying to find?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

\[(2)(10)(300)=(2)(\sqrt{51})x'\]

OpenStudy (anonymous):

so 300 mph isn't ds/dt?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i got 420.08 mph :)